If you drop an aspirin from rest and it takes 0.35 seconds to hit the table, how high above the table was the aspirin when it was released?

a.
1.72 m

b.
1.2 m

c.
0.6 m

d.
3.43 m
1
Expert's answer
2020-02-10T09:30:31-0500

"S=v_0 \cdot t+a \cdot t^2",

where "S" is a displacement, "v_0" - initial speed, "t" - time, "a" - acceleration (in considered case acceleration equals to the gravitational acceleration "a=g \approx 9.8 \; m/s^2" ).

"S=0 \cdot 0.35+9.8 \cdot 0.35^2 \approx 1.2 \; m"


Answer: b. 1.2 m.